OpenStudy (anonymous):

Show ALL work for the following problem |4x + 7| - 2 = 29

pooja195 (pooja195):

Any ideas on what to do first?

OpenStudy (anonymous):

add 2 on both sides?

pooja195 (pooja195):

Yep what are we left with?

rebeccaxhawaii (rebeccaxhawaii):

|absolute value signs| whatevers in it positive but since they are already positive they stay the same

OpenStudy (anonymous):

4x+7=31

pooja195 (pooja195):

Good 4x+7=31 absolute value's have 2 solutions one postitive and one negative soo 4x+7=-31 4x+7=31 solve both for x

OpenStudy (anonymous):

x=6 x=-9.5

pooja195 (pooja195):

Perfect :)

OpenStudy (anonymous):

thank you @pooja195